OpenStudy (anonymous):

Which expressions have a value of –15? Choose all answers that are correct. A. –2 – (–13) B. 7 – 22 C. –24 – (–9) D. –4 – 19

pooja195 (pooja195):

Solve each one and see which equals -15

OpenStudy (anonymous):

do u know how to solve them?

OpenStudy (anonymous):

kinda 1 sec

OpenStudy (anonymous):

a and idk

OpenStudy (anonymous):

y did u choose A?

OpenStudy (anonymous):

its supostu be -15

OpenStudy (anonymous):

yeah so what r the answers

OpenStudy (anonymous):

A isnt one of them

OpenStudy (anonymous):

the answers are B and C... the equal to -15

OpenStudy (anonymous):

@cupcakeohla

pooja195 (pooja195):

Lets solve them one by one –2 – (–13) two neg signs together make a positive –2 +13 This can also be written as 13-2=?

pooja195 (pooja195):

13 MINUS 2= ?

OpenStudy (anonymous):

11

pooja195 (pooja195):

Good so we know A is incorrect

pooja195 (pooja195):

7 – 22 This can be written as -22+7=?

OpenStudy (anonymous):

15

pooja195 (pooja195):

Well it would be -15 :P But good so we know B is an answer Next –24 – (–9) This can be written as -24+9 =?

OpenStudy (anonymous):

-15

pooja195 (pooja195):

Good so C is also an answer –4 – 19 this can also be written as -19-4=?

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

so ther all right

pooja195 (pooja195):

When we subtract a negative it increases (negative wise) -19-4= -23 not -15.. B and C are the only right ones

OpenStudy (anonymous):

ok thx

pooja195 (pooja195):

yw

OpenStudy (anonymous):

could u help me again

pooja195 (pooja195):

Sure post it as a new question please.

OpenStudy (mathstudent55):

@pooja195 rocks!

pooja195 (pooja195):

lol thanks <3 xD

OpenStudy (skullpatrol):

+1

pooja195 (pooja195):

:3

TheSmartOne (thesmartone):

+1

jaynator495 (jaynator495):

+3